OpenStudy (anonymous):

The second term in a geometric sequence is 20. The fourth term in the same sequence is 45/4, or 11.25. What is the common ratio?

OpenStudy (anonymous):

Ok..see!! so you know the formula for a term in a geometric sequence an=a×rn−1 a is the 1st term and r is the common ratio. so in your question you know 2nd term 20=ar2−1...or.....20=ar and the 4th term 11.25=ar4−1....or....11.25=ar3 the 4th term can be written as ar times r^2 since r * r^2 = r^3 so you can substitute then 2nd term 11.25=ar×r2....or........11.25=20×r2 now you just need to solve for r. -----------> and r=0.75 i guess :O Hope it helps a bit...!

OpenStudy (anonymous):

is that 11.25=20 times r squared or 20x times r squared?

OpenStudy (anonymous):

20 times r squared.

OpenStudy (anonymous):

THANK YOU! :)

OpenStudy (anonymous):

yeah np..Anytime :)